Passes through (-6,-4) and (-1,9)
Answer by mukhopadhyay(490)

Use formula (y-y1)/(x-x1) = (y2-y1)/(x2-x1)
=> (y+4)/(x+6) = (9+4)/(-1+6)
=> (y+4)/(x+6) = 13/5
=>5y+20 = 13x+78
=> 13x-5y = -58
Answer: Desirable equation is 13x-5y = -58